OpenStudy (anonymous):

If -4(r+6)=-63,what does r=?

OpenStudy (anonymous):

39/4

jimthompson5910 (jim_thompson5910):

\[\large -4(r+6)=-63\] \[\large -4r-24=-63\] \[\large -4r=-63+24\] \[\large -4r=-39\] \[\large r=\frac{39}{4}\]